AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 358     Entry time: Tue Mar 1 23:30:04 2022
Author: David, George 
Subject: Current status 

00:30
Pressure in S1 slightly decreased from before but still high enough to measure data.
One new count at 20MeV

Target density at S1: 2e-8 - 2.7e-8
Pressures OK (one small spike just before 00:00 in YR08)
System Checks OK
White rabbit fail diff: 49
Data write speed: ~1.4Mb/s
Beam intensity at injection: ~7.7e7
Cryo Temp Setpoint: 110K

(Beam intensity in ring dropped as I was writing which is why there is no rutherford peak on exRate image)
 

Attachment 1: 2022-03-02_00-39-47xytotal.png  195 kB  | Hide | Hide all
2022-03-02_00-39-47xytotal.png
Attachment 2: 2022-03-02_00-39-20exRate.png  155 kB  | Hide | Hide all
2022-03-02_00-39-20exRate.png
Attachment 3: 2022-03-02_00-38-44extotal.png  167 kB  | Hide | Hide all
2022-03-02_00-38-44extotal.png
ELOG V3.1.4-unknown